OpenStudy (anonymous):

Look at the sequence below. What is the next term in the sequence? [51 , 47 , 43 , 39 , 35] A. 34 B. 32 C. 31 D. 29

OpenStudy (anonymous):

okay umm is there a pattern?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

51-47= 4 47-43=4 so...35-4= 31 woo!

OpenStudy (anonymous):

thanks :]

OpenStudy (anonymous):

Anytime.